# Concept-Aware LoRA for Domain-Aligned Segmentation Dataset Generation

### v1

@@ E.3 Additional Qualitative Results
Figure 16: Qualitative comparison between the original LoRA and our Style CA-LoRA for generating the “person” class to increase the label proportion. While the original LoRA can generate the “person” class, it is limited in producing informative samples beyond the training set, with generated images often resembling those from the training set. In contrast, the Style CA-LoRA generates diverse scenes for the “person” class, as it exclusively learns the style from the source dataset.
@@
F Class-Specific Dataset Generation
F.1 Class-wise Segmentation Performance
@@
Class-wise Performance Improvements (Fig. 15)
@@
Urban-scene segmentation has distinct challenges, including class imbalance and co-occurrence issues [25], making class-wise analysis particularly important. We present the class-wise IoU improvements in Fig. 15. As shown in Fig. 15 (a), our proposed dataset generation approach proves especially effective for rare classes such as “bus”, “fence”, and “bicycle”. However, the generated dataset often fails to improve performance in certain classes, such as “person” and “rider”. As illustrated in Fig. 15 (b), this degradation is primarily due to the insufficient number of generated samples for the “person” class. Since the synthetic dataset is generated randomly, disparities in label proportions can occur. To mitigate this, we propose a simple yet effective technique to increase the proportion of the target class.
Method Training Dataset Total Iterations IoU mIoU
# Real # Generated Car Bus Motorcycle
Baseline 9 90K 84.02 12.51 16.11 41.83
For a fair comparison, we fine-tune the baseline for 30K iterations using the following datasets.
Baseline (FT) 9 120K 83.98 13.37 15.96 42.00
Ours 9 500 120K 85.03 30.59 15.26 44.12
Ours (+ Diverse Cars) 9 900 120K 85.34 32.48 17.77 44.95
@@
Table 11: In-domain segmentation performance of datasets incorporating the diverse cars dataset. Incorporating the diverse cars dataset especially improved performance for vehicle classes such as “car”, “bus”, and “motorcycle”, leading to overall performance improvements. Since we generated an additional 400 image-label pairs (100 images per vehicle type), the total number of the generated samples is 900.
Desired Concept Proportion of T2I Model Layers for Concept-Aware Fine-tuning
0% (Pretrained) 1% 2% 3% 5% 10% 100% (Original LoRA)
Style 42.82 43.77 44.13 43.94 43.81 43.36 42.97
Viewpoint 43.13 43.01 42.37 42.94 42.52
Table 12: Hyperparameter search for the proportion of selected layers in few-shot segmentation (Cityscapes, 0.3%). Style CA-LoRA achieves the best performance with a 2% selection of layers.
Segmentation Dataset Generation Focused on a Specific Class (Figs. 15 and 16)
@@
As detailed in Section 3.4 and Table 13, we generated the dataset using the prompt “photorealistic first-person urban street view with [Class names]”, where the class names were extracted from the label map of the training set by retrieving the names of all classes present in the label map. While the synthesized text prompt partially reflects the label proportions of the training set, it does not strictly enforce these proportions. As a result, the proposed generated dataset may exhibit misaligned label proportions, as illustrated in Fig. 15 (b).
@@
To address this issue, we propose a class-specific generation approach that manually increases the target class by modifying the generation prompts. Specifically, we generated an additional 500 samples using the prompt “photorealistic first-person urban street view with people” to increase the proportion of the “person” class. [^6] We also experimented with “photorealistic first-person urban street view with person”, but using “people” as the test prompt proved to be more effective in increasing the label proportion for the “person” class. Since we selectively fine-tuned the LoRA to learn only the style from Cityscapes, it enables effective manipulation using the text prompt, which the original LoRA cannot achieve, as demonstrated in Fig. 16. As illustrated in Fig. 15 (b), this approach successfully increased the proportion of the “person” class and mitigated its performance degradation. Furthermore, as shown in Fig. 15 (a), this adjustment led to additional performance improvements, increasing the average IoU from 44.12 to 44.59.
F.2 Generating Datasets with Diverse Class Names
@@
Since the Style CA-LoRA selectively fine-tuned only the style from the in-domain Cityscapes dataset, it retains its generalization ability for text prompts such as objects. Leveraging this capability, we aim to generate a broader variety of images using more diverse class names beyond those provided in the dataset. In this experiment, we refined the prompts for generating images previously created with the simple class name “car” by subdividing them into “sedan car”, “SUV car”, “convertible car”, and “hatchback car”, as shown in Fig. 17. As illustrated in the figure, while the original LoRA fine-tuned text-to-image generation model struggles to produce diverse styles of cars, our approach reliably generates a wide variety of cars that align with the test prompts.
@@
G Implementation Details
G.1 Label Generator Architecture
Since DatasetDM is built on top of Stable Diffusion v1.5 [41], we simply adjust the feature dimensions in the projection layers to accommodate Stable Diffusion XL [37]. The detailed label generator architecture is illustrated in Fig. 18. However, if all feature maps and cross-attention maps are used, the total number of channels increases significantly, leading to an unmanageable number of parameters in the projection layers during concatenation and projection. In summary, the feature maps are extracted from the last feature block at each resolution of the up-sampling blocks, while cross-attention maps are sampled at equal intervals (every 7 blocks) from the total 36 up-sampling blocks (i.e., 1st, 8th, … 29th, 36th).
Furthermore, as shown in Fig. 18, Stable Diffusion XL has only three resolution levels, compared to four resolution levels of the Stable Diffusion v1.5 architecture in the original DatasetDM. In the Mask2Former structure, feature maps from the pixel decoder, excluding the largest resolution, are fed into the transformer decoder. While the original design used three-resolution feature maps, only two were utilized in this case. Thus, while DatasetDM provides three-resolution feature maps three times for 9 transformer decoder blocks, we provide two-resolution feature maps five times, leading to a total of 10 transformer decoder blocks. Importantly, to ensure a fair comparison, the reported scores for DatasetDM were obtained using a re-implemented version based on SDXL with the same modifications.
G.2 Detailed Architecture of CA-LoRA
Overview (Fig. 19 (a))
The fundamental group of weights used to measure concept sensitivity is the linear projection layer. We selectively adapt the pretrained weights layer by layer within the projection layers. Practically, the projection layers in the multi-head attention block are integrated across heads, we split them to structurally distinguish the weights for each projection. Consequently, we attach LoRA layers projection-wise, as shown in Fig. 19.
Projection-wise CA-LoRA (Fig. 19 (b))
$\Delta W_{\text{OUT}}=B\begin{bmatrix}&&&\\ A_{1}&A_{2}&\cdots&A_{h}\\ &&&\end{bmatrix},$ (10)
The projection-wise LoRA is represented in Fig. 19 and Algorithm 4.
G.3 Hyperparameters and Pseudocode
Hyperparameters (Tables 13, LABEL:, 14, LABEL:, 15, LABEL: and 16)
Pseudocode
We also provide PyTorch-like pseudocode [36] for key algorithms to effectively support reproducibility. Concept sensitivity (Algorithms 1 and 2) The concept sensitivity algorithm is demonstrated in Algorithm 1. Conducting concept sensitivity requires several helper functions, as shown in Algorithm 2. CA-LoRA (Algorithms 3 and 4) The CA-LoRA algorithm is divided into two parts: the forward function and the declaration function. The forward pass of CA-LoRA is presented in Algorithm 3, while the declaration function, along with the selected layers, is illustrated in Algorithm 4.
Figure 20: Qualitative results of Style- and Viewpoint CA-LoRA according to the layer proportions (1%, 2%, 3%, 5%, and 10%). While the Style and Viewpoint CA-LoRA effectively disentangle with the small proportions of the selected layers, it has been entangled according to the increased proportion of the selected layers. This flexibility allows for manual adjustment of the extent to which other concepts are learned, depending on the specific problem settings or datasets.
Figure 21: Qualitative examples of the generated image-label pairs for in-domain and domain generalization settings. Style CA-LoRA effectively generates a Cityscapes-style dataset. Viewpoint CA-LoRA can control the weather condition of the generated images with corresponding label maps since it selectively learns the Cityscapes-viewpoint.
Figure 22: Additional examples of our generated datasets for fully supervised and domain generalization scenarios.
@@ G.3 Hyperparameters and Pseudocode
Test prompt augmentation (In-domain) “… with [Class names]”77footnotemark: 7
@@
unet = unet.requres_grad_(True)
@@
sensitivity.append(grads_concept / grads_diff)
sensitivity_avg = average_gradients(sensitivity) # Algorithm 2
Algorithm 1 PyTorch-like Pseudocode of Concept Sensitivity
@@
proj = getattr_recursive(attn_module, proj_name.split(’.))
@@
Algorithm 2 PyTorch-like Helper Functions for Concept Sensitivity

### v2

@@ Appendix E Class-Specific Dataset Generation
Figure 17: Qualitative comparison between the original LoRA and CA-LoRA for generating “person” samples. While the original LoRA produces limited variations resembling training images, CA-LoRA generates more diverse “person” scenes by learning only the style from the source dataset.
@@
Figure 19: Additional qualitative results
E.1 Class-wise Segmentation Performance
@@
Class-wise Performance Improvements (Figure 16)
@@
Urban-scene segmentation has distinct challenges, including class imbalance and co-occurrence issues [25], making class-wise analysis particularly important. We present the class-wise IoU improvements in Figure 16. As shown in Figure 16 (a), our proposed dataset generation approach proves especially effective for rare classes such as “bus”, “fence”, and “bicycle”. However, the generated dataset often fails to improve performance in certain classes, such as “person” and “rider”. As illustrated in Figure 16 (b), this degradation is primarily due to the insufficient number of generated samples for the “person” class. Since the synthetic dataset is generated randomly, disparities in label proportions can occur. To mitigate this, we propose a simple yet effective technique to increase the proportion of the target class.
@@
Table 11: In-domain segmentation performance of datasets incorporating the diverse cars dataset. Incorporating the diverse cars dataset especially improved performance for vehicle classes such as “car”, “bus”, and “motorcycle”, leading to overall performance improvements.
Method IoU mIoU
Car Bus Motorcycle
Baseline 84.02 12.51 16.11 41.83
Ours 85.03 30.59 15.26 44.12
Ours (+ Diverse Cars) 85.34 32.48 17.77 44.95
Segmentation Dataset Generation Focused on a Specific Class (Figures 16 and 17)
@@
As detailed in Section 3.4 and Table 12, we generated the dataset using the prompt “photorealistic first-person urban street view with [Class names]”, where the class names were extracted from the label map of the training set by retrieving the names of all classes present in the label map. While the synthesized text prompt partially reflects the label proportions of the training set, it does not strictly enforce these proportions. As a result, the proposed generated dataset may exhibit misaligned label proportions, as illustrated in Figure 16 (b).
@@
To address this issue, we propose a class-specific generation approach that manually increases the target class by modifying the generation prompts. Specifically, we generated an additional 500 samples using the prompt “photorealistic first-person urban street view with people” to increase the proportion of the “person” class. [^3] We also experimented with “photorealistic first-person urban street view with person”, but using “people” as the test prompt proved to be more effective in increasing the label proportion for the “person” class. Since we selectively fine-tuned the LoRA to learn only the style from Cityscapes, it enables effective manipulation using the text prompt, which the original LoRA cannot achieve, as demonstrated in Figure 17. As illustrated in Figure 16 (b), this approach successfully increased the proportion of the “person” class and mitigated its performance degradation. Furthermore, as shown in Figure 16 (a), this adjustment led to additional performance improvements, increasing the average IoU from 44.12 to 44.59.
E.2 Generating Datasets with Diverse Class Names
@@
Since the Style CA-LoRA selectively fine-tuned only the style from the in-domain Cityscapes dataset, it retains its generalization ability for text prompts such as objects. Leveraging this capability, we aim to generate a broader variety of images using more diverse class names beyond those provided in the dataset. In this experiment, we refined the prompts for generating images previously created with the simple class name “car” by subdividing them into “sedan car”, “SUV car”, “convertible car”, and “hatchback car”, as shown in Figure 18. As illustrated in the figure, while the original LoRA fine-tuned text-to-image generation model struggles to produce diverse styles of cars, our approach reliably generates a wide variety of cars that align with the test prompts.
@@
Test prompt augmentation (In-domain) “… with [Class names]”44footnotemark: 4
@@
unet = unet.requires_grad_(True)
@@
awareness.append(grads_concept / grads_diff)
awareness_avg = average_gradients(awareness) # Algorithm 2
Algorithm 1 PyTorch-like Pseudocode of Concept Awareness
@@
proj = getattr_recursive(attn_module, proj_name.split(’.’))
@@
Algorithm 2 PyTorch-like Helper Functions for Concept Awareness